OpenStudy (anonymous):

a sum of twice a number and 4 is equal to the sum of the number and 6

Directrix (directrix):

@ritzy Let N = the number Twice the number = 2N Sum of the number and 4 = N + 4 Now, translate from words to symbols

Directrix (directrix):

a sum of twice a number and 4 is equal to the sum of the number and 6 2N + 4 = N + 6 So, solve the equation: 2N + 4 = N + 6 for N. Post what you get and we'll compare answers, okay? @ritzy